OpenStudy (anonymous):

Choose the correct simplification of the expression (2x2y6z5)(5x4y5z3). 10 x8y30z15 7 x6y11z8 7x8y30z15 10x6y11z8

OpenStudy (anonymous):

There's no correct choice I can't tell which one is x or multiply, or exponent :( (2x2y6z5)(5x4y5z3) Multiply by terms appropriately 2x * 5x = 10x² 2y * 4y = 8y² 6z * 5z = 30z^8

OpenStudy (anonymous):

that's what i said

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

LOL YOU JUST COPIED FROM ANOTHER SITE NICE HELP. https://answers.yahoo.com/question/index?qid=20111209185803AA54w8I